OpenStudy (anonymous):

let f(x)= (3/x) and g(x) = (4/x^2)-1. find f(g(x)). simplify your answer.

OpenStudy (anonymous):

I get \[3/((4/x^2)-1)\] but i do not know how to simplify it

OpenStudy (zzr0ck3r):

f((4/x^2)-1) = 3/((4/x^2)-1)) = 3/((4-x^2)/x^2) = 3x^2/(4-x^2)

OpenStudy (anonymous):

thanks